### v1

# thanks: These two authors contributed equallythanks: These two authors contributed equally

Abstract
We show that the three commonly employed approaches that define the same intrinsic linear anomalous Hall response actually lead to different results for intrinsic nonlinear transport. The difference arises from an intrinsic anomalous distribution. It originates from scattering, but its value is completely independent of scattering, because it represents the local equilibration of electron wave packets with field corrected energy. As a manifestation, we find that under ac driving, the intrinsic contributions in rectified component and in double-frequency component exhibit distinct frequency dependence, which can be probed in experiment. Using first-principles calculations, we estimate the signals that can be probed in antiferromagnetic CuMnAs.
Intrinsic responses, which depend only on band structures of materials, are of crucial importance in condensed matter physics, since they can be evaluated accurately with first-principles techniques, offer benchmarks for experiments, and probe interesting geometric band properties. A prominent example is anomalous Hall effect [1, 2], whose intrinsic contribution manifests the Berry curvature of band structures [3, 4]. In that context, there have been three different approaches to define what is the intrinsic response in diffusive transport. (1) The first approach is to consider the so-called “clean limit”, without any scattering. For example, in the semiclassical theory [5, 2], this means one directly takes the equilibrium distribution $f_{0}$ in the calculation of response. (2) The second is to define the intrinsic response as the extrapolation of the ac Hall response from the high-frequency regime ($\omega\tau\gg 1$, where $\tau$ is the scattering time) to the dc limit, with $1/\tau\rightarrow 0$ faster than $\omega\rightarrow 0$ [1]. (3) The third one is an operational and intuitive definition, namely, one fully considers scattering and defines intrinsic contribution as the scattering independent component of the total response. For anomalous Hall transport, the three approaches are considered equivalent, in the sense that they yield the same intrinsic response.
Recently, there has been tremendous interest in nonlinear Hall effect, for which the charge current response $j\sim E^{2}$ is quadratic in the applied $E$ field [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Under broken time reversal symmetry, this effect also allows an intrinsic contribution [6, 29, 40]. Despite intensive studies, the nature of this intrinsic nonlinear transport has not been thoroughly understood. Particularly, it remains unclear whether the three approaches above still define the same intrinsic response in the nonlinear regime. If not, which one will be a more proper definition?
In this work, we address the above fundamental problem. We find that distinct from linear response, the three approaches actually lead to three different results for second order nonlinear transport (see Table. 1). This difference originates from the field correction to the electron wave packet energy, which must be included for nonlinear response as it is quadratic in $E$. Owing to this energy shift, in the presence of disorder scattering, electrons acquire a correction in distribution function, the intrinsic anomalous distribution, which turns out to be completely independent of scattering. Physically, the corrected distribution can be naturally interpreted as the equilibrium state for an ensemble of electron wave packets with the corrected energy. The establishment of this new equilibrium requires scattering, but the resulting distribution does not depend on scattering. It follows that approach 1 must differ from approach 3, since it does not consider any scattering effect; meanwhile, approach 2 partially captures the above contribution and leads to a result sitting in-between. We further investigate the situation under ac driving. Interestingly, we find that even within the intraband transport domain, the intrinsic rectified response $j^{(0)}$ and the intrinsic double-frequency response $j^{(2\omega)}$ have different amplitudes in approach 3. In addition, $j^{(2\omega)}$ also exhibits different values in low-frequency and high-frequency regimes. Based on these insights, we argue that approach 3 fits into a natural and consistent physical picture and also connects closely with experimental measurement, so it appears to be a more appropriate definition. We also discuss possible ways to test some of our predictions by nonlinear measurement on antiferromagnetic CuMnAs.
Table 1: Different results of second order intrinsic nonlinear transport under the three definitions of intrinsic contribution which are equivalent for linear response.
Component without counting scattering $\ \omega\tau\gg 1$ and $\omega\rightarrow 0$  with scattering, expressions independent of disorder
dc limit    Eq. (6)  $\frac{1}{2}$(Eq. (6)+Eq. (7))    Eq. (7)
ac rectified    Eq. (6)   --    Eq. (7)
ac double-frequency    Eq. (6)   --    $\omega\tau\gg 1$, Eq. (6);  $\omega\tau\ll 1$, Eq. (7)
Preliminaries. In the semiclassical theory, the charge current density is expressed as (we set $e=\hbar=1$)
$\bm{j}=-\sum_{l}f_{l}\bm{v}_{l},$ (1)
where $l\equiv(n,\bm{k})$ is a composition of band and crystal momentum labels, and $\bm{v}_{l}$ is the velocity of the electron wave packet. Up to second order in applied uniform $E$ field, the scattering-independent part of velocity is given by [6, 42]
$\bm{v}_{l}=\partial_{\bm{k}}\tilde{\varepsilon}_{l}+\bm{E}\times\tilde{\bm{\Omega}}_{l}.$ (2)
Here,
$\tilde{\varepsilon}_{l}=\varepsilon_{l}+\delta\varepsilon_{l}$ (3)
is the field-corrected wave packet energy: Besides the unperturbed band energy $\varepsilon_{l}$, there is, as mentioned, a second order correction [43, 44, 37]
$\delta\varepsilon_{l}=-\frac{1}{2}G_{ab}^{l}E_{a}E_{b},$ (4)
where summation over repeated Cartesian indices $a$ and $b$ is implied, and $G_{ab}^{l}$ is known as Berry-connection polarizability (BCP) tensor [24, 43]. It is a gauge-invariant intrinsic band property, given by [6]
$G_{ab}^{l}=2\text{Re}\sum_{n^{\prime}\neq n}\frac{V_{a}^{nn^{\prime}}V_{b}^{n^{\prime}n}}{(\varepsilon_{n}-\varepsilon_{n^{\prime}})^{3}},$ (5)
where $V^{nn^{\prime}}$ is the interband velocity matrix element, and the quantities on the right hand side are evaluated at the same $k$ point. In Eq. (2), the $k$-space Berry curvature $\tilde{\bm{\Omega}}_{l}$ includes the first-order field correction, i.e., $\tilde{\bm{\Omega}}=\bm{\Omega}+\delta\bm{\Omega}$, where $\bm{\Omega}$ is for unperturbed bands, and interestingly, the correction $\delta\Omega_{a}=\epsilon_{abc}\partial_{k_{b}}G_{cd}E_{d}$ ($\epsilon_{abc}$ is the totally antisymmetric tensor) is also determined by BCP [6, 23, 24].
To obtain the intrinsic nonlinear current, in approach 1, one simply put $f_{l}=f_{0}(\varepsilon_{l})$, substitute Eq. (2) into (1), and collect terms of $E^{2}$ order. The result is
$\bm{j}^{\mathrm{cl}}=-\sum_{l}f_{0}(\varepsilon_{l})\Big{(}\partial_{\bm{k}}\delta\varepsilon_{l}+\bm{E}\times\delta\bm{\Omega}_{l}\Big{)}.$ (6)
One may already spot an ambiguity in the treatment above, namely, since we are effectively dealing with a bunch of field-dressed wave packets with energies $\tilde{\varepsilon}_{l}$, it seems more appealing to use their equilibrium distribution $f_{0}(\tilde{\varepsilon}_{l})$ in Eq. (1) for the intrinsic current. By doing so and performing the expansion $f_{0}(\tilde{\varepsilon})=f_{0}({\varepsilon})+f_{0}^{\prime}({\varepsilon})\delta\varepsilon$, we obtain a different and rather simple intrinsic nonlinear current:
$\bm{j}^{\mathrm{in}}=-\sum_{l}f_{0}(\varepsilon_{l})(\bm{E}\times\delta\bm{\Omega}_{l}),$ (7)
which is a purely Hall (hence dissipationless) response. The two currents differ by
$\bm{j}^{\mathrm{in}}-\bm{j}^{\mathrm{cl}}=-\sum_{l}g_{l}\partial_{\bm{k}}\varepsilon_{l},$ (8)
where
$g_{l}=f_{0}(\tilde{\varepsilon}_{l})-f_{0}({\varepsilon}_{l})$ (9)
is the difference between the two distributions, which we refer to as the intrinsic anomalous distribution.
In fact, $\bm{j}^{\mathrm{in}}$ in Eq. (7) has been obtained in several pioneering works on this effect [6, 23, 24]. We shall uncover the scattering origin of its difference from $\bm{j}^{cl}$ and see below that it in fact corresponds to the intrinsic response defined by approach 3.
The role of scattering. Now, the question is: Where does the intrinsic anomalous distribution $g_{l}$ come from? The original system in the absence of driving field is in the equilibrium state described by $f_{0}(\varepsilon_{l})$. When one gradually turns on the $E$ field, the electron wave packets are dressed, with their energies changed to $\tilde{\varepsilon}_{l}$. Clearly, $f_{0}(\varepsilon_{l})$ makes no longer an equilibrium state for particles with these renormalized energies, so it should relax towards an (new) equilibrium corresponding to $f_{0}(\tilde{\varepsilon}_{l})$. This is the part of the distribution function in addition to the non-equilibrium part due to the drift of Fermi surface by field and other extrinsic effects. Evidently, the relaxation from $f_{0}(\varepsilon_{l})$ to $f_{0}(\tilde{\varepsilon}_{l})$ must be through scattering.
The picture above can be justified by examining the Boltzmann equation, which describes the kinetics of electron wave packets. Assuming uniform system and field, it reads $\partial_{t}f_{l}-\bm{E}\cdot\partial_{\bm{k}}{f_{l}}=\mathcal{I}\{f_{l}\}$, where the right hand side is the collision integral, acting as a linear operator on $f_{l}$. Note that the $E$-field driving term here produces the non-equilibrium part $\sim E\tau$ in the distribution function, which does not contribute to the intrinsic current. Therefore, to study intrinsic response, we may drop this term and focus on the solution of
$\partial_{t}f_{l}=\mathcal{I}\{f_{l}\}.$ (10)
Note that this equation is describing wave packets with energies $\tilde{\varepsilon}_{l}$. In other words, the scattering processes in $\mathcal{I}$ are defined for these dressed wave pacekts. For example, in the case of one-electron elastic scattering, we have [45]
$\mathcal{I}\{f_{l}\}=-\sum_{l^{\prime}}(\omega_{l^{\prime}l}f_{l}-\omega_{ll^{\prime}}f_{l^{\prime}}),$ (11)
with the scattering rate $\omega_{l^{\prime}l}=2\pi|T_{l^{\prime}l}|^{2}\delta(\tilde{\varepsilon}_{l^{\prime}}-\tilde{\varepsilon}_{l}),$ where $T$ is the scattering $T$-matrix [2]. Then, it is clear that $f_{0}(\tilde{\varepsilon}_{l})$ is always a static solution of (10) (an explicit treatment of this simple case can be found in the Supplemental Material [46]). In contrast, $f_{0}({\varepsilon}_{l})$ is not a static solution: If imposed as an initial distribution, it should evolve by Eq. (10) and relax through scattering towards the static solution $f_{0}(\tilde{\varepsilon}_{l})$ (In this process, one may assume that chemical potential and temperature of the system are fixed by environment).
The above analysis is general, regardless of the details of the system as well as scattering. It follows that there always exists the intrinsic anomalous distribution $g_{l}$ in the dc limit, which originates from scattering but its value is independent of scattering. Thus, in approach 3, the intrinsic nonlinear current should be evaluated with $f_{0}(\tilde{\varepsilon}_{l})$ instead of $f_{0}({\varepsilon}_{l})$, and the result indeed takes the form of Eq. (7).
Frequency dependence. It remains to find the intrinsic response defined by approach 2. In this approach, we need to consider the frequency dependence due to ac driving. Let us take $\tilde{E}_{a}(t)=E_{a}\cos\omega t=\text{Re}(E_{a}e^{i\omega t})$. At $E^{2}$ order, the current response contains two frequency components: $j_{a}=\text{Re}(j_{a}^{(0)}+j_{a}^{(2\omega)}e^{i2\omega t})$, a zero-frequency (rectified) component $j^{(0)}$ and a double-frequency component $j^{(2\omega)}$. Both components may contain intrinsic contributions.
The distribution function can also be decomposed into frequency components, with
$f_{l}=f_{0}(\varepsilon_{l})+\text{Re}\Big{(}g^{(0)}_{l}+g^{(\omega)}_{l}e^{i\omega t}+g^{(2\omega)}_{l}e^{i2\omega t}\Big{)}.$ (12)
It is known that the $g^{(\omega)}$ component contributes only to extrinsic terms [20], so we focus on $g^{(0)}$ and $g^{(2\omega)}$.
From the analysis of Eq. (10) just now, one can readily see that the zero-frequency component $g^{(0)}$ always contains the following intrinsic anomalous distribution
$g^{(0)}_{l}=f_{0}^{\prime}(\varepsilon_{l})\delta\varepsilon^{(0)}_{l}=-\frac{1}{4}f_{0}^{\prime}(\varepsilon_{l})G_{ab}^{l}E_{a}E_{b},$ (13)
where $\delta\varepsilon^{(0)}_{l}$ denotes the zero-frequency component of the energy correction. This intrinsic anomalous distribution is independent of frequency. Here, we assumed $\omega<\Delta$ with $\Delta$ the size of local gap at Fermi surface, such that the intraband transport dominates the response.
Meanwhile, for the double-frequency component $g^{(2\omega)}$, its value depends on frequency and differs in two regimes. In the low-frequency regime $\omega\ll\tau^{-1}$, the left hand side of (10) is negligible, so the result takes the same form as the zero-frequency component:
$g_{l}^{(2\omega)}\approx f_{0}^{\prime}(\varepsilon_{l})\delta\varepsilon^{(2\omega)}_{l}=g^{(0)}_{l}.$ (14)
On the other hand, in the high-frequency regime $\tau^{-1}\ll\omega<\Delta$, the right hand side of (10) becomes negligible, and we obtain
$g_{l}^{(2\omega)}\approx 0.$ (15)
This can be readily understood, since at such high frequency, the oscillating electrons simply do not “see” the scatterers. For frequency between the two regimes, $g_{l}^{(2\omega)}$ should exhibit explicit $\omega$ dependence, and crossover from one regime to the other. In the Supplemental Material [46], we show that for the case of elastic scattering with relaxation time approximation for the Boltzmann equation, the crossover takes the form of
$\text{Re}\Big{(}g_{l}^{(2\omega)}e^{i2\omega t}\Big{)}=g^{(0)}_{l}\frac{\cos 2\omega t+2\omega\tau\sin 2\omega t}{1+(2\omega\tau)^{2}}.$ (16)
Inputting these results into Eq. (1), one finds that, in the sense of approach 3, for the low-frequency regime $\omega\ll\tau^{-1}$, the ac intrinsic nonlinear current is given by
$j^{(0)}_{a}=j^{(2\omega)}_{a}=\frac{1}{2}j^{\mathrm{in}}_{a},$ (17)
where $j^{\mathrm{in}}_{a}$ is given by expression (7) with $E$ the amplitude of ac $E$ field. As for the high-frequency regime $\tau^{-1}\ll\omega<\Delta$, the result is
$j^{(0)}_{a}=\frac{1}{2}j^{\mathrm{in}}_{a},\qquad j^{(2\omega)}_{a}=\frac{1}{2}j^{\mathrm{cl}}_{a}.$ (18)
Therefore, the double-frequency intrinsic nonlinear response exhibits difference between low and high frequency regimes, whereas the rectified response does not. This frequency dependence is a unique feature of nonlinear response that has not been revealed before. It has important experimental implications that will be discussed below.
Now, if we use approach 2 to define the dc intrinsic nonlinear current, i.e., taking Eq. (18) and extrapolating to dc limit, the result is given by $(1/2)(\bm{j}^{\mathrm{cl}}+\bm{j}^{\mathrm{in}})$. Thus, the three approaches give three different results for the intrinsic nonlinear response. All these results are summarized in Table 1.
Discussion. We have revealed that the three approaches, which define the same intrinsic linear anomalous Hall response, produce different results when applied to the nonlinear effect. The question is which one is more reasonable. Their difference comes from the intrinsic anomalous distribution function $g_{l}$. As we have shown, although $g_{l}$ has its root in scattering, its value is completely independent of scattering. It arises from the natural equilibration of electrons under a local shift of energy. It follows that in dc measurement, one can never separate $g_{l}$ from $f_{0}(\varepsilon_{l})$; in other words, they always come together in the form of $f_{0}(\tilde{\varepsilon}_{l})$ in producing the response. Thus, for better comparison between theory and experiment, approach 3 seems to be the more proper definition.
It is worth noting that the nature of this intrinsic anomalous distribution contribution is different from the so-called side jump contribution in anomalous Hall effect [47, 2]. In early literature, it is often stated that the side jump contribution is independent of scattering. However, this conclusion is not valid in general: it has been shown that the side jump contribution depends on the type of disorders and generally varies in the presence of more than one scattering sources [48, 49]. In contrast, the intrinsic anomalous distribution is completely independent of these details, so it can be regarded as an intrinsic response.
The result from approach 3, such as Eq. (7), is also physically appealing. In the picture of semiclassical theory, we are dealing with dynamics of a bunch of electron wave packets. The external fields not only supply driving force, they also dress the wave packets, leading to their renormalized energy $\tilde{\varepsilon}_{l}$ and Berry curvature $\tilde{\Omega}$. Then, in a consistent treatment, it is natural that the response is formulated in terms of these dressed wave packets with their renormalized properties. In this sense, the use of $f_{0}(\tilde{\varepsilon}_{l})$ for intrinsic response is a physically consistent choice. In the same spirit, for a consistent treatment, the scattering amplitude $|T_{l^{\prime}l}|^{2}$ should be defined with respect to the dressed wave packets, which has been shown in Ref. [50].
The same logic applies to the case under applied magnetic field. It is well known that to linear order in $B$, the wave packet energy acquires a correction $\delta\varepsilon_{l}=-\bm{m}_{l}\cdot\bm{B}$, where $\bm{m}_{l}$ is the spin plus orbital magnetic moment of the wave packet [51, 5]. Then, the study of magneto-responses should also include the intrinsic anomalous distribution due to this correction $\delta\varepsilon_{l}$. Notably, it has been shown that failing to account for this intrinsic anomalous distribution will lead to the unphysical result that a static $B$ field alone can drive a transport current [52, 53, 54]. This is a strong evidence showing the necessity to include the intrinsic anomalous distribution for a consistent theory.
It is worth noting that implicit behind the picture is the separation between two time scales. One is the time scale for local equilibration among wave packets, which drives the distribution towards $f_{0}(\tilde{\varepsilon}_{l})$. It is associated with the relaxation of the local energy shift $\delta\varepsilon_{l}$ (This is also the process that establishes $f_{0}(\varepsilon_{l})$ in the absence of field). The other is the time scale for the macroscopic transport, which might be regarded as a relaxation of the electric potential energy. The latter is nonlocal in the sense that the relaxation requires the transport of electron across a macroscopic distance (through the sample). Clearly, in a diffusive system, the former time scale (on the order of $\tau$) is much shorter than the latter. This justifies the use of $f_{0}(\tilde{\varepsilon}_{l})$ as a starting point for studying transport process.
Figure 1: (a) Crystal structure of orthorhombic CuMnAs. The local moment direction in the antiferromagnetic ground state is marked. (b) Calculated band structure. (c) Calculated nonlinear response $\chi^{(2\omega)}_{xxx}$ versus chemical potential $\mu$. Temperature is set at 303030~{}30K. (d) Distribution of BCP dipole (the summand in Eq. (19) in the $k_{y}=0$ plane in momentum space. The dashed line marks the nodal loop in the band structure around $X$ point, which is slightly gapped under spin-orbit coupling.
Finally, we note that the frequency dependence of $j^{(2\omega)}$ under ac driving [Eq. (17) versus (18)] is a result of this new understanding, which can be probed in experiment. Experimentally, one typically applies an ac driving at frequency $\omega$ and measures the double-frequency response using lock-in technique [10, 11]. From our results, one can see that at low frequencies $\omega\ll\tau^{-1}$, the $2\omega$ response is purely Hall (dissipationless). If $\tau$ is sufficiently large that one can tune $\omega$ into the regime $\tau^{-1}\ll\omega<\Delta$, then there will also appear a non-Hall (dissipative) component in the transverse double-frequency response. This can be distinguished in experiment by the sum-frequency generation method developed by Gao et al. [29]. Meanwhile, in the high-frequency regime, there also emerges a double-frequency component in longitudinal response. The corresponding nonlinear conductivity is
$\chi_{aaa}^{(2\omega)}=-\sum_{l}V_{a}^{ll}G_{aa}^{l}f^{\prime}_{0}(\varepsilon_{l})/2.$ (19)
To have an estimation of this quantity, we consider the example of orthorhombic CuMnAs, which is an antiferromagnet with Néel temperature $T_{\mathrm{N}}\approx 300$ K [55, 56, 57, 58, 59, 60, 61]. Previous experiments [58, 59] reported that the Néel vector is aligned along the $b$ axis [Fig. 1(a)]. The magnetic band structure obtained from first-principles calculations (details in [46]) is shown in Fig. 1(b), which agrees with previous calculations [58]. This material belongs to the $m^{\prime}mm$ magnetic point group. The symmetry-allowed nonlinear longitudinal response is $\chi_{xxx}$. The result calculated from Eq. (19) is shown in Fig. 1(c). Interestingly, we find that the sizable $\chi_{xxx}$ has a large contribution from a nodal loop feature near the Fermi level [see Fig. 1(d)]. At $T=30~{}$K, the value of $\chi_{xxx}$ can reach $\sim 2.3~{}$mA///V2. From previous experiment [59], $\tau\sim 1$ ps at 30 K, and from the band structure, $\Delta\sim 20$ meV ($30$ THz). Hence, this scattering-independent nonlinear response $\chi_{xxx}$ can be observed in the frequency window $\mathrm{0.5~{}THz}\ll\omega<\mathrm{30~{}THz}$. This implies that orthorhombic CuMnAs is a promising candidate for terahertz isolation [62, 63, 64] of intrinsic nonlinear transport.